OpenStudy (ja1):

What is the simplified form of the expression:

OpenStudy (ja1):

\[8\sqrt{v^2} - \sqrt{v^4} - \sqrt{9v^2} + \sqrt{16v^4}\]

OpenStudy (ja1):

@jim_thompson5910 @zepdrix @ganeshie8 @dan815 @some_someone

OpenStudy (ja1):

@amistre64 sorry to bother you for such a simple problem but no one comes <_<

OpenStudy (amistre64):

you have all even exponents; dividing by 2 gets rid of the sqrts

OpenStudy (amistre64):

\[8\sqrt{v^2} - \sqrt{v^4} - \sqrt{9v^2} + \sqrt{16v^4}\] \[8\sqrt{v^2} - \sqrt{v^4} - 3\sqrt{v^2} + 4\sqrt{v^4}\] \[8v^{2/2} - v^{4/2} - 3v^{2/2} + 4v^{4/2}\] \[8v - v^{2} - 3v + 4v^{2}\]

OpenStudy (ja1):

Ohh ok I did that but I did some things wrong but I get it now

OpenStudy (ja1):

Thanks a bunch saved me a lot of time xD

OpenStudy (ja1):

That's a neat way to do it to, I haven't learned it like that yet but it will come in handy and seems very quick too.